### v1

# Inferring activity from the flow field around active colloidal particles using deep learning

Abstract
Active colloidal particles create flow around them due to non-equilibrium process on their surfaces. In this paper, we infer the activity of such colloidal particles from the flow field created by them via deep learning. We first explain our method for one active particle, inferring the $2s$ mode (or the stresslet) and the $3t$ mode (or the source dipole) from the flow field data, along with the position and orientation of the particle. We then apply the method to a system of many active particles. We find excellent agreements between the predictions and the true values of activity. Our method presents a principled way to predict arbitrary activity from the flow field created by active particles.
1 Introduction
Active colloidal particles – such as microorganisms (Brennen & Winet, 1977; Goldstein, 2015), and synthetic microswimmers (Ebbens & Howse, 2010; Thutupalli et al., 2018) – are known to organise themselves in emergent structures via fluid-mediated interactions. An active particle can create flow around it even if it is not moving due to non-equilibrium process on its surface, which could be of biological origin (Brennen & Winet, 1977) or of phoretic origin (Ebbens & Howse, 2010). Classifying the irreducible components of flow produced by active particles provides a handle through which we can tune inter-agent interactions to design higher order (hierarchical) assemblies of active particles which display a desired self-organization and tunable emergent dynamics.
A topic of recent interest is the application of tools from machine learning to bear upon questions of inference and design principles in active matter systems (Cichos et al., 2020; Boffi & Vanden-Eijnden, 2024; Tsang et al., 2020; Supekar et al., 2023; Boffi & Vanden-Eijnden, 2024). Convolutional neural networks (CNNs) have been used for performing Particle Image Velocimetry (Rabault et al., 2017), particle tracking(Newby et al., 2018) and feature extraction(Hannel et al., 2018). To the best of our knowledge, a work on estimating activity from the flow field around active particles has not been reported. To this end, we use deep learning techniques given fluid flow around particles to infer the irreducible components of activity.
The remainder of the paper is organized as follows. In section 2, we describe a method to compute irreducible components of flow around active particles. The methodology utilising deep learning to learn activity from fluid flow around active particles is given in section 3. We apply the method to construct irreducible components of flow around an active particle in section 4. We extend this method to infer flow around many particles in section 5. Finally, we conclude in section 6 by summarizing our main results and suggesting directions for future investigations.
2 Irreducible components of flow around active particles
In this section, we consider the problem of computing irreducible components of flow around active colloidal particles in a Stokesian fluid of viscosity $\eta$. At low Reynolds number, as appropriate to active colloidal particles, the fluid flow $\mathbf{v}$ satisfies the Stokes equation (Lauga, 2020)
$-\mathbf{\nabla}p+\eta\nabla^{2}\mathbf{v}$ $=-\mathbf{f},\qquad\mathbf{\nabla}\cdot\mathbf{v}=0.$ (1)
Here $p$ is the fluid pressure and $\mathbf{f}$ is the force density. In appendix A, we show that the flow around active particles can be written in terms of irreducible components $\mathbf{v}^{l\sigma}(\mathbf{k})$. The sum of the irreducible components in the Fourier space is:
${\mathbf{v}}^{M}(\mathbf{k})=\sum_{l=1}^{\infty}\sum_{\sigma\in\{s,a,t\}}\mathbf{v}^{l\sigma}(\mathbf{k}).$ (2)
As described in the appendix A, the flow $\mathbf{v}^{l\sigma}(\mathbf{r})$ decays as $r^{-l}$. Here $\sigma$ can take three values for each $l$: $\sigma=s$, $\sigma=a$, and $\sigma=t$, which correspond to the three irreducible mode allowed for each $l$. See appendix A for details. The terms $\mathbf{v}^{1s}$ is the flow due to a net external force, while $\mathbf{v}^{2a}$ is the flow due to a net external torque. Since an active particle has no external force or torque on it, the two leading irreducible components of exterior flow created by active particles are:
$\mathbf{v}^{2s}(\mathbf{k})$ $=\,\mathbf{G}(\mathbf{k})\cdot\mathbf{f}^{2s},\qquad\mathbf{f}^{2s}=i\mathbf{k}\cdot\left[M\mathbf{Q}^{(2s)}\right]$ (3)
$\mathbf{v}^{3t}(\mathbf{k})$ $=\mathbf{G}(\mathbf{k})\cdot\mathbf{f}^{3t},\qquad\mathbf{f}^{3t}=-k^{2}M\,\mathbf{Q}^{(3t)}.$ (4)
The above two irreducible components of flow - the $2s$ mode (or the stresslet (Lauga, 2020)) and the $3t$ mode (or the source dipole (Lauga, 2020)) - is known to fit flow around a wide range of active particles (Lauga, 2020; Marmol et al., 2024; Goldstein, 2015; Ghose & Adhikari, 2014; Delmotte et al., 2015; Drescher et al., 2010, 2011; Baruah et al., 2025; Thutupalli et al., 2018; Krishnamurthy & Prakash, 2023; Guasto et al., 2010). The leading fluid flow due to a force-free and torque-free active particles is due to the two modes: $\mathbf{Q}^{(2s)}$ and $\mathbf{Q}^{(3t)}$. In addition, we can use uniaxial parameterization of the single layer density, which is true for a wide range of microorganisms having symmetry about a given axis. For such a scenario, with $\mathbf{p}$ being the orientation of the particle, we choose:
$\mathbf{Q}^{(2s)}=s_{0}\left(\mathbf{p}\mathbf{p}-\frac{1}{n_{d}}\mathbf{I}\right),\qquad\mathbf{Q}^{(3t)}=d_{0}\,\mathbf{p}$ (5)
Here $n_{d}$ is the space dimension and $\mathbf{I}$ is the identity tensor, while $s_{0}$ and $d_{0}$ are constants. We note that the constants $s_{0}$ and $d_{0}$ are related (Singh et al., 2015) to the $B_{1}$ and $B_{2}$ modes of the squirmer model (Lighthill, 1952; Blake, 1971; Pedley, 2016; Lauga, 2020). The flow field due to these two modes is present in Fig. 1.
[…]
For reasons explained above, we truncate the expression of the fluid flow in Eq.(2) to the leading modes $2s$ (stresslet) and $3t$ (source dipole). For this choice, the expression for the net force density in the fluid, defined in Eq.(1), is written as the sum:
$\mathbf{f}(\mathbf{r})=\mathbf{f}^{2s}+\mathbf{f}^{3t}=-\mathbf{\nabla}M(\mathbf{r})\cdot\mathbf{Q}^{(2s)}-{\nabla}^{2}M(\mathbf{r})\,\mathbf{Q}^{(3t)}.$ (6)
See appendix A for details. Here, $M$ is a Gaussian mollifier whose mean is at the centre of the particle and the variance equals $2b^{2}/\pi$ (Maxey & Patel, 2001), where $b$ is the radius of the particle. In the next section, we present a deep learning method to infer the individual components of active force density (and also the positions and the orientations of the particles) given a fluid flow data from the flow field around single and many active particles.
3 Methodology
Finding the velocity field solution to the Stokes equation with a given force density is discussed in the previous section. In addition, the velocity field may also be obtained from experimental systems using methods such as Particle Image Velocimetry (PIV)(Adrian et al., 1991; Willert & Gharib, 1991). On the other hand, the inverse problem of finding the force density (and their irreducible components) from a given flow field data is non-trivial. This leads us to a neural-network-based learning approach for achieving the force density and corresponding irreducible modes using the universal function approximator capacity of neural networks (Hornik et al., 1989).
To effectively approximate the inverse function using deep learning, we begin by constructing a sufficiently large and diverse dataset. This dataset is partitioned into training, validation, and test subsets. The model is trained on the training set to learn the mapping from velocity to force density. Throughout training, the validation set is used to monitor accuracy. Further the model is evaluated on the independent test set to verify model’s performance on unseen data.
We consider flow with two possibilities, a single particle in a grid and 4 particles in a grid which can be generalized to $N$ particles. We generate two datasets of 4000 realizations corresponding to systems with different particle counts on a $128\times 128$ domain. The grid is divided into four equal sub-regions, and one particle is randomly placed within each to ensure non-overlapping configurations in case of more than one particle.
The $i$th particle, whose center is located at $\mathbf{c}^{(i)}=(x^{(i)},y^{(i)})$, is assigned an orientation vector $\mathbf{p}^{(i)}=(\cos\theta^{(i)},\sin\theta^{(i)})$, where $\theta^{(i)}$ is an angle made by the orientation vector with the $\hat{x}$-axis. The angle $\theta^{(i)}$ is selected from a discretized set of angles with $1^{\circ}$ resolution. Discretization of orientation space facilitates improved learning by the neural network. Each particle is assigned randomized stresslet strength $s_{0}\in[800,1200]$ and source dipole strength $d_{0}\in[12000,16000]$. The data is partitioned into training, validation, and test sets comprising 2800, 600, and 600 samples, respectively.
The neural network used in this work is a fully convolutional architecture designed to approximate the inverse problem of Eq.(1). The input is a two-channel ($x$ and $y$ components of the fluid velocity $\mathbf{v}$) tensor $\mathbf{v}\in\mathbb{R}^{H\times W\times 2}$, representing the force field on a two-dimensional grid of height $H$ and width $W$. The network output is $\tilde{\mathbf{f}}\in\mathbb{R}^{H\times W\times 2}$, which approximates the force field for the given velocity field. Each convolutional layer ${(l)}$ applies a localized linear transformation followed by a nonlinear activation function (Goodfellow et al., 2016). The network consists of a composition of such layers as shown in fig. 1. The operation at layer $l$ is given by:
$z^{(l+1)}=g(W^{(l)}*z^{(l)}+b^{(l)})=h^{(l)}(z^{(l)})\qquad g(z)=\max(0,z)$ (7)
where $z^{(l)}\in\mathbb{R}^{H\times W\times C_{l}}$ is the input to the layer, $W^{(l)}\in\mathbb{R}^{C_{l+1}\times C_{l}\times k\times k}$ is a learnable convolutional kernel of size $k\times k$, and $b^{(l)}\in\mathbb{R}^{C_{l+1}}$ is a bias vector. The symbol * denotes the two-dimensional convolution operation. $g$ is a nonlinear activation, ReLU, defined in Eqn. 7. ReLU is preferred over the sigmoid activation function because it helps prevent the saturation of artificial neurons(Glorot et al., 2011). To improve generalization and reduce overfitting, dropout is applied after each activation with a dropout probability of $0.2$. The neural network is trained with Mean Squared Error(MSE) as loss function and Adam(Kingma & Ba, 2014) as optimizer.
Figure 2: Flowchart of the method for the predictions of the model for a single particle: The velocity field for inference is shown along with a schematic of the neural network. The neural network predicts the force density field (here only magnitude is shown for clarity) from input velocity field. The velocity field obtained from predicted force is reconstructed with good accuracy. See Algorithm 1 for details of the method. The flow streamlines have been overlaid on a pseudo color plot of the speed.
As we have explained in the further sections the network is able to compute force density for both single particle and multiple particle case accurately. We can also infer about the contribution of stresslet and source dipole in the flow field. The cases for one and many particles are analyzed differently, but the neural network for inversion is same for both the cases.
4 Inferring one particle flow
Active and self-propelled particles in the absence of applied forces or torques, with uniaxial parametrization, can be characterized by the flow generated. The stresslet $\mathbf{Q}^{(2s)}$ generates flows that neither translate nor rotate the particle. The translational effect is generated by the source dipole $\mathbf{Q}^{(3t)}$. The sign of stresslet strength $s_{0}$ with respect to source dipole strength $d_{0}$ determines how the active particle acts. It behaves as ’puller’, forcing the fluid inward along the swimming direction and repel fluid from their sides,if $s_{0}/d_{0}>0$ and as ’pusher’, force fluid outward along their swimming direction and draw fluid into the sides of their bodies, if $s_{0}/d_{0}<0$.
We consider the case of a single particle in a $128\times 128$ grid, which is a slice from a large grid with periodic boundaries. The velocity field of the grid is fed into the neural network as input. The network computes the force density $\mathbf{\tilde{f}}=(\tilde{f}_{x},\tilde{f}_{y})$ by approximating the inverse problem accurately as shown in 2. The orientation $(\mathbf{p})$ of the stresslet is computed by taking local average of predicted force density in the neighborhood where the magnitude of predicted force density is maximum. The position of the particle, strength of stresslet and the source dipole strengths are estimated by minimizing the squared error of true force density and estimated force density with gradient descent. The algorithm for estimating these parameters is given in Algorithm 1.
Algorithm 1 Particle Detection and Parameter Estimation for Many Particles
0:  Force density $\tilde{f}_{x}$, $\tilde{f}_{y}$; threshold ratio $\tau$; particle size $b$; minimum distance $d_{\text{min}}$
1:  Compute force density magnitude: $\tilde{f}_{\text{mag}}(x,y)\leftarrow\sqrt{\tilde{f}_{x}^{2}(x,y)+\tilde{f}_{y}^{2}(x,y)}$
2:  Obtain global maximum: $\tilde{f}^{*}_{\text{mag}}\leftarrow\displaystyle\max_{x,y}\tilde{f}_{\text{mag}}(x,y)$
3:  Detect peaks:
$\mathcal{P}\leftarrow\left\{(x,y):(x,y)\notin\mathbf{c}^{(j)}\pm d_{\text{min}}\ \text{and}\ \tilde{f}_{\text{mag}}(x,y)>\tau\cdot\tilde{f}^{*}_{\text{mag}}\right\}$
4:  Detect number of particles $N\leftarrow||\mathcal{P}||$ and compute center of mass $\{\mathbf{c}^{(i)}\}_{i=1}^{N}$ for each region of $2d_{min}\times 2d_{min}$ centered around peak.
5:  for each $\mathbf{c}^{(i)}$ do
6:     Initialize $f_{0}^{(i)},d_{0}^{(i)}$
7:     Estimate local direction: $\mathbf{p}^{(i)}\leftarrow(\bar{\tilde{f}}_{x},\bar{\tilde{f}}_{y})\quad\text{in neighborhood of }\mathbf{c}^{(i)}$
8:     Initialize best loss: $\mathcal{L}_{\min}\leftarrow\infty$
9:     for $\mathbf{c}^{(i)}\leftarrow\mathbf{c}^{(i)}+\Delta\mathbf{c}$ do
10:        Define estimated force field using gaussian envelope$(M)$:
$\mathbf{\hat{f}}=s_{0}\nabla M(\mathbf{c}^{(i)})\cdot\left(\mathbf{p}^{(i)}\mathbf{p}^{(i)}-\frac{\mathbf{I}}{2}\right)-d_{0}\nabla^{2}M\left(\mathbf{c}^{(i)}\right)\mathbf{p}^{(i)}$
11:        Minimize loss:
$\mathcal{L}(s_{0},d_{0})=\sum_{\Omega}\left\|\mathbf{\tilde{f}}-\mathbf{\hat{f}}\right\|^{2}$
12:        Update parameters if $\mathcal{L}<\mathcal{L}_{\min}$
13:     end for
14:     Store $(s_{0}^{(i)},d_{0}^{(i)},\mathbf{p}^{(i)},\mathbf{c}^{(i)})$
15:  end for
16:  return  $\{(f_{0}^{(i)},d_{0}^{(i)},\mathbf{p}^{(i)},\mathbf{c}^{(i)})\}_{i=1}^{N}$
The prediction of neural network gives a very accurate representation of force density when compared to the ground truth as seen in figure 2. We reconstruct the velocity field using the parameters obtained from the predicted density using Algorithm 1. The predicted velocity field represents the flow with a good accuracy and we are able to decompose the flow into $2s$ and $3t$ modes.
The predictions of the stresslet and source dipole strength give very close result with respect to the ground truth having $3.26\%$ and $5.70\%$ error respectively averaged over 600 test cases unseen to the neural network. We have predicted the parameters using the neural networks for particles with arbitrary positions and orientations and reconstructed the velocity field(Figure 3).
Figure 3: True (top row) and predicted (bottom row) velocity fields due to different values of stresslet strength $(s_{0})$, source dipole strength $(d_{0})$, orientation $(\mathbf{p})$ and particle position $(\mathbf{c})$. Values have been rounded to the nearest integer for clarity. We show four distinct examples for illustration. The flow velocity streamlines have been overlaid on a pseudo color plot of the flow speed.
5 Inferring many particle flow
We extend our model further to test with flow due to $N$ particles. Here we consider the case of $N=4$. In an arbitrary velocity field we first detect the number of particles in the box detecting number of locations where force density magnitude is concentrated over some small local area using scikit-ndimage (Van der Walt et al., 2014). We ensure there is sufficient distance between two different regions associated with particles while detecting. The approximate particle position can be guessed as center of mass (force density magnitude) of that region. Now each of these regions can be thought of a single particle system and the estimation methods of single particle can be applied. The process is explained in Algorithm 1.
The Neural Network is able to predict the force density effectively as seen from the comparison with ground truth force density in Fig. 4. From the predicted density we get the parameters by Algorithm 1 and use it to reconstruct the velocity field. Thus, our method using deep learning to extract force density and then fitting that data to the known model for the fluid flow allows prediction in a system of many particles.
Figure 4: Predictions of the model for multiple particles in a grid. The flow velocity streamlines have been overlaid on a pseudo color plot of the flow speed. Magnitude of the force density is also shown. Our inference method using deep learning predicts location and orientation of the particles as well as the force density in the fluid due to them.
6 Summary and Conclusions
In this paper, we have studied the inverse problem of computing activity of colloidal particles from the specified flow field data around them by combining artificial intelligence and physics-based methods. Such a method is particularly attractive since PIV data from experiments can provide us with flow field data. Using deep learning, we predict activity from the flow field data created by active particles. We achieve this by inverting the flow field data to obtain force density in the fluid using CNN (Convolutional Neural Networks). The reconstructed flow field from the predicted force density is in excellent agreement with the true flow field. Finally, using a new inference algorithm, we infer the position and orientation of active particles along with irreducible components of the velocity (stresslet and source dipole).
In this paper, we have only focused on two modes - $2s$ and $3t$ widely used to model flow around active particles (Lauga, 2020; Marmol et al., 2024; Goldstein, 2015; Ghose & Adhikari, 2014; Delmotte et al., 2015; Drescher et al., 2010, 2011; Baruah et al., 2025; Thutupalli et al., 2018; Krishnamurthy & Prakash, 2023; Guasto et al., 2010). We would like to note that the method presented here is not limited to these two modes and it can be applied to arbitrary modes of active slip around active particles for a given flow data. Indeed, our method presents a principled way to compute irreducible components of activity around each particle. In this paper, we have used deterministic fluid velocities around active particles in two-dimension to infer their activity. Our method can also be applied for time-varying activity given a series of snapshots of velocity fields to predict how these inferred parameters change with time. The method presented here can also be be extended to cases when there is noise in the fluid velocity by denoising and reconstruction methods (Weng et al., 2001; Wang et al., 2022). Our work may also be extended to include other higher modes and consider data from three-dimensional systems. Comparison of different architectures to find the most optimal one for this application is another exciting direction for future work. For generality, we have used a simple loss function in our CNN architecture. Extending the current work to a physics informed neural network with a more problem-specific choice of loss function suggest another exciting avenue for future work.
Appendix A Solving the Stokes equation to compute flow around active particles
The flow around a finite sized rigid sphere of radius $b$ in a fluid of viscosity $\eta$ with single layer density $\mathbf{q}$ is given by the boundary integral expression (Pozrikidis, 1992):
[…]
The above integrals can be solved by expanding the boundary fields in TSH (tensorial spherical harmonics) (Hess, 2015). The TSH are defined as $Y_{\alpha_{1}\dots\alpha_{l}}^{(l)}(\hat{\mathbf{b}})=(-1)^{l}\,b^{(l+1)}\,\nabla_{\alpha_{1}}\dots\nabla_{\alpha_{l}}\frac{1}{b}$. The expansion of the single layer density $\mathbf{q}$ in TSH is:
$\mathbf{q}(\mathbf{b})$ $=\sum_{l=1}^{\infty}w_{l}\,\mathbf{Q}^{(l)}\cdot\mathbf{Y}^{(l-1)}(\mathbf{\hat{\mathbf{b}}}),\qquad\mathbf{Q}^{(l)}=\tilde{w}_{l}\int\mathbf{q}\,\mathbf{Y}^{(l-1)}(\mathbf{\hat{\mathbf{b}}}),$ (9a)
Here, $\mathbf{b}$ is the radius vector of active colloidal sphere, while the $l-$dependent constants are: $w_{l}=\frac{1}{(l-1)!(2l-3)!!}$ and ${w}_{l}=\frac{2l-1}{4\pi b^{2}}$. Here, we have defined $M!!=M(M-2)(M-4)\dots$, while $\mathbf{\Delta}^{(l)}$ is a tensor of rank $2l$ projecting the symmetric irreducible part of a tensor of rank $l$ (Hess, 2015). In the above, and throughout the paper, a maximal contraction of two tensors is denoted by a dot product.
The boundary integral of Eq.(8) can be solved exactly as follows: (a) expanding the boundary fields in TSH, (b) performing a Taylor expansion of kernels (the Green’s function) about the center of the particles, and (c) using the orthogonality of TSH (Singh et al., 2015; Singh, 2018). The final result is the following summation:
$\mathbf{v}(\mathbf{r})=-\sum_{l=1}^{\infty}\left(1+\frac{b^{2}}{4l+2}\nabla^{2}\right)\mathbf{\nabla}^{(l-1)}\mathbf{G}(\mathbf{r})\cdot\mathbf{Q}^{(l)},$ (10)
Thus, the flow due to $l$th mode decays as $r^{-l}$, since $\mathbf{G}(\mathbf{r})\propto r^{-1}$.
The equation (1) in the Fourier space is:
$\mathbf{v}(\mathbf{k})$ $=\mathbf{G}(\mathbf{k})\cdot\mathbf{f}(\mathbf{k}),\quad\quad\mathbf{G}(\mathbf{k})=\frac{1}{\eta k^{2}}\left({\mathbf{I}}-\frac{\mathbf{k}\mathbf{k}}{k^{2}}\right).$ (11)
We define Fourier transform of a function $\varphi(\mathbf{r})$ and Dirac delta function $\delta(\mathbf{r})$ as
[…]
Using equations (11) and (10), we can identify an effective force density:
$\mathbf{f}(\mathbf{k})=-\sum_{l=0}^{\infty}\left(1-\frac{b^{2}}{4l+2}k^{2}\right)(i\mathbf{k})^{(l-1)}\cdot\mathbf{Q}^{(l)}$ (13)
that produces the flow in Eq. (10). Using the definition of Dirac delta function, the above force density can be written in the coordinate space as:
$\mathbf{f}(\mathbf{r})=-\sum_{l=1}^{\infty}\left(1+\frac{b^{2}}{4l+2}\nabla^{2}\right)\mathbf{\nabla}^{l-1}\delta(\mathbf{r})\cdot\mathbf{Q}^{(l)}.$ (14)
that satisfies the Stokes equation. Since our interest is in the external flow, we are free for this purpose to replace the $\delta$ function in the above expression with a mollifier $M$ which produces the same exterior flow (Maxey & Patel, 2001; Delmotte et al., 2015; Singh, 2018). We choose a Gaussian mollifier whose mean is at the center of the particle and the variance equals $2b^{2}/\pi$ (Maxey & Patel, 2001). Using Eq.(11), the expression of the flow in Fourier space is
${\mathbf{v}}^{M}(\mathbf{k})$ $=-\sum_{l=1}^{\infty}\left(1-\frac{b^{2}}{4l+2}k^{2}\right)(-i\mathbf{\mathbf{k}})^{l-1}\,M(\mathbf{k})\mathbf{G}(\mathbf{k})\cdot\mathbf{Q}^{(l)}$ (15)
The coefficients of single layer density $\mathbf{Q}^{(l)}$ are $l$-th rank tensors, symmetric irreducible in their last $l-1$ indices. The reducible tensors $\mathbf{Q}^{(l)}$ can be written in irreducible forms as: $\mathbf{Q}^{(l\sigma)}=\mathbf{P}^{(l\sigma)}\cdot\mathbf{Q}^{(l)}$ Turk et al. (2024). In the above expression, the operator $\mathbf{P}^{(l\sigma)}$ projects symmetric irreducible ($\mathbf{P}^{(ls)}=\mathbf{\Delta}^{(l)}$), antisymmetric ($\mathbf{P}^{(la)}=\mathbf{\Delta}^{(l-1)}\mathbf{\mathbf{\varepsilon}}$), and trace ($\mathbf{P}^{(lt)}=\mathbf{\delta}$) parts of a reducible tensor of rank-$l$ (Hess, 2015). Here $\mathbf{\Delta}^{(l)}$ is a tensor of rank $2l$ projecting the symmetry irreducible part of a tensor of rank $l$, $\mathbf{\varepsilon}$ is the Levi-Civita tensor and $\mathbf{\delta}$ is the Kronecker delta (Hess, 2015). $\mathbf{Q}^{(l\sigma)}$ are irreducible tensors of rank $l$, $l-1$ and $l-2$ for $\sigma=s,a,$ and $t$ respectively. The final expression of the flow in terms of irreducible components can be written as per Eq.(2) of the main text. For simplicity and less error in computation the symmetric-dipole ($2s$) can also be written as two forces acting along $\mathbf{p}$ and $-\mathbf{p}$ each $\Delta a$ length apart from the center of particle. This formulation of $2s$ doesn’t keep the traceless property, but that is taken care in the flow field by the incompressibility condition.